2. Suppose that the banks of the nation of Atlantis have a reserve ratio of 20%, and there is a new deposit of $500,

and that the people of Atlantis have no desire to hold currency.

a) Complete the following table.

Round           Deposits            Required Reserves          Excess Reserves               Loans

1                         $500                    $100                                             $400                    $400

2                        $400

3

4

5

6

7

8

9

10

Total after 10 rounds

b) What is the money multiplier?
Answer #1

Solution:-

(A).

Round

Deposits

Required Reserves

Excess Reserves

Loans

A

B = A * 20%

C = A - B

1

$500

$100

$400

$400

2

$400

$80

$320

$320

3

$320

$64

$256

$256

4

$256

$51.2

$204.8

$204.8

5

$204.8

$40.96

$163.84

$163.84

6

$163.84

$32.77

$131.07

$131.07

7

$131.07

$26.21

$104.86

$104.86

8

$104.86

$20.97

$83.89

$83.89

9

$83.89

$16.78

$67.11

$67.11

10

$67.11

$13.42

$53.69

$53.69

Total

2231.56

446.31

1785.25

1785.25

(B). Money Multiplier = 1 / Required Reserve

                                         = 1 / 0.20

                                         = 5